# The dark side of the forces: assessing non-conservative force models for atomistic machine learning

### v3

@@ 2.1 Interatomic potentials and atomistic simulations
An interatomic potential $V$ describes the potential energy between the NN atoms of a molecule or material as a function of their positions $\{\mathbf{r}_{i}\}_{i=1}^{N}$ and chemical nature $\{a_{i}\}_{i=1}^{N}$:
@@
$\mathbf{f}_{j}=-{\partial V(\{\mathbf{r}_{i},a_{i}\}_{i=1}^{N})}\mathbin{\big/\penalty 50}{\partial\mathbf{r}_{j}}\,.$ (2)
@@
MD in the NVE ensemble. The number of atoms NN, volume $V$ and total energy $E$ (kinetic plus potential) are kept constant, describing the behavior of an isolated system. Usually, this is accomplished by using the velocity Verlet integrator (Verlet, 1967) with a short time step $\Delta t$.
@@ 2.4 Breaking physical constraints
The conservative property for forces in Eq. (5) is satisfied if and only if there exists a function (usually named the potential energy) of which the forces are the spatial derivatives, see Eq. (2). Therefore, a machine learning model that calculates interatomic forces as derivatives of the potential energy according to Eq. (2) will be trivially conservative. The required derivatives, the gradient of the scalar $V$, can be obtained efficiently (i.e., with the same asymptotic computational cost as the energy prediction) with automatic differentiation (Griewank & Walther, 2008). Nevertheless, differentiation incurs a computational overhead, typically $2\mathrm{-}3\times$ for inference and $3\times$ for training; the exact theoretical factors are discussed further in Appendix B. This overhead can be avoided by directly predicting forces during the forward pass. However, by removing the relationship between forces and energy defined in Eq. (2), such models do not enforce energy conservation. The possibility of performing a direct, non-conservative, evaluation of the forces was realized early (Li et al., 2015), but used only sparsely until recently, when it was applied to equivariant graph neural networks such as GemNet (Gasteiger et al., 2021) and several more recent universal machine learning interatomic potentials including ORB (Neumann et al., 2024) and Equiformer (Liao et al., 2023), which are trained on large datasets and promise broad applicability for chemical modeling. Some of these architectures also perform direct evaluation of stresses, which leads to the violation of conservation of enthalpy (see App. G.3).
@@ 3.1 Measuring non-conservative behavior
[…]
where $\mathbf{J}_{\mathrm{anti}}=(\mathbf{J}-\mathbf{J}^{\top})/2$. $\lambda$ then defines a metric going from 0 for conservative forces to 11 for forces that have no conservative component. $\lambda$ can also be computed only for entries associated with an atom $i$, or a pair of atoms ijij, providing a finer-grained assessment of the violation of Eq. 6, as seen in Figure 1.
@@ 4.4 Constant-energy molecular dynamics
Given that non-conservative models lack a well-defined conserved quantity by construction, we rely on indirect measurements of the sampled ensemble. We consider in particular the kinetic temperature $T=2K/(3Nk_{\mathrm{B}})$, where NN is the number of particles considered. This is just a rescaling of the kinetic energy $K$; its ensemble average should correspond to the target temperature for NVT trajectories (300 K in these tests), and to a value in its vicinity for NVE trajectories initialized from a thermally equilibrated configuration. As a sensitive indicator of the dynamical behavior of the system, we compute the Fourier transform of the velocity-velocity correlation function, $\hat{c}_{vv}(\omega)$. Its peaks are closely related to the density of vibrational modes and to infrared and Raman spectra, and its $\omega\rightarrow 0$ limit is proportional to the diffusion coefficient.
@@ 4.8 Non-conservative forces as accelerators
In simulations, one can then use the conservative forces of such a hybrid model for validation, error monitoring and correction, and the direct forces for faster inference. A good example is to use multiple time-stepping (MTS) techniques (Tuckerman et al., 1992) for molecular dynamics, where the non-conservative forces are used to integrate the equations of motion, and the conservative forces are applied every $M$ steps as a correction. This reduces the theoretical overhead of a conservative trajectory from a factor of $F\approx 2$ to one of 1+(F1)/M1+(F-1)/M. The results using this technique in Table 2, Figure 2 and in Appendix I are essentially indistinguishable from fully conservative ones, using $M=8$, which leads to a small, approximately 2020% slowdown compared to a direct-force, non-conservative trajectory. Appendix I contains further explanation of the MTS technique, as well as more detailed results for MTS simulations using models trained on the water and OC20 datasets. The technique can be also successfully used for constant-pressure simulations, as shown in Section G.3.
@@ Appendix B Theoretical computational cost
However, in the case of backward force evaluation, operations where $x$ is an internal representation and $y$ is a weight can save the $\partial V/\partial y$ calculation. In a simple multi-layer perceptron, where all linear layers correspond to this type of computation, this would yield a backward propagation that is roughly as expensive as the forward pass. This is not the case in transformers, as the attention mechanism involves comparatively expensive operations where both $x$ and $y$ are internal representations, and one can expect the backward propagation of gradients to be somewhere between $1\times$ and $2\times$ the cost of the forward evaluation.
@@ G.2 NVT
[…]
@@ G.6 Generalized Langevin equations
[…]
@@ Appendix I Multiple time stepping
where $\mathbf{q}$ and $\mathbf{p}$ indicate the vectors containing positions and momenta of all particles, and $\mathbf{m}$ the masses. Forces have to be recomputed if they are needed after a position update. In the limit of ideal time scale separation, the resulting dynamics (and the sampled ensemble for constant-temperature simulations) is consistent with $V_{\text{slow}}$, at a much-reduced cost: Letting $F$ denote the slowdown of “slow” over “fast” forces, and assuming that an evaluation of “slow” forces also yields the “fast” ones, MTS reduces the theoretical cost of well-behaved molecular dynamics from $F$ to 1+(F1)/M1+(F-1)/M. Choosing $M$ therefore requires a trade-off between accuracy and speed. Morrone et al. (2011) discusses the failure modes of MTS at high $M$, and shows how to use thermostats for mitigation.
@@
ORB, MACE-MP-0, Equiformer, SevenNet calculations were performed using the publicly available repositories, respectively https://github.com/orbital-materials/orb-models/ (orb-models 0.4.0), https://github.com/ACEsuit/mace (mace-torch 0.3.8), https://github.com/FAIR-Chem/fairchem (fairchem-core 1.3.0), and https://github.com/MDIL-SNU/SevenNet (sevenn $0.1$).

### v4

@@ 2.1 Interatomic potentials and atomistic simulations
An interatomic potential $V$ describes the potential energy between the $N$ atoms of a molecule or material as a function of their positions $\{\mathbf{r}_{i}\}_{i=1}^{N}$ and chemical nature $\{a_{i}\}_{i=1}^{N}$:
@@
$\mathbf{f}_{j}=-{\partial V(\{\mathbf{r}_{i},a_{i}\}_{i=1}^{N})}\mathbin{\big{/}}{\partial\mathbf{r}_{j}}\,.$ (2)
@@
MD in the NVE ensemble. The number of atoms $N$, volume $V$ and total energy $E$ (kinetic plus potential) are kept constant, describing the behavior of an isolated system. Usually, this is accomplished by using the velocity Verlet integrator (Verlet, 1967) with a short time step $\Delta t$.
@@ 2.4 Breaking physical constraints
[…]
@@ 3.1 Measuring non-conservative behavior
$\lambda=\frac{||\mathbf{J}_{\mathrm{anti}}||_{\text{F}}}{||\mathbf{J}||_{\text{F}}}\,,$ (7)
where $\mathbf{J}_{\mathrm{anti}}=(\mathbf{J}-\mathbf{J}^{\top})/2$. $\lambda$ then defines a metric going from 00 for conservative forces to $1$ for forces that have no conservative component. $\lambda$ can also be computed only for entries associated with an atom $i$, or a pair of atoms $ij$, providing a finer-grained assessment of the violation of Eq. 6, as seen in Figure 1.
@@ 4.4 Constant-energy molecular dynamics
Given that non-conservative models lack a well-defined conserved quantity by construction, we rely on indirect measurements of the sampled ensemble. We consider in particular the kinetic temperature $T=2K/(3Nk_{\mathrm{B}})$, where $N$ is the number of particles considered. This is just a rescaling of the kinetic energy $K$; its ensemble average should correspond to the target temperature for NVT trajectories (300 K in these tests), and to a value in its vicinity for NVE trajectories initialized from a thermally equilibrated configuration. As a sensitive indicator of the dynamical behavior of the system, we compute the Fourier transform of the velocity-velocity correlation function, $\hat{c}_{vv}(\omega)$. Its peaks are closely related to the density of vibrational modes and to infrared and Raman spectra, and its $\omega\rightarrow 0$ limit is proportional to the diffusion coefficient.
@@ 4.8 Non-conservative forces as accelerators
In simulations, one can then use the conservative forces of such a hybrid model for validation, error monitoring and correction, and the direct forces for faster inference. A good example is to use multiple time-stepping (MTS) techniques (Tuckerman et al., 1992) for molecular dynamics, where the non-conservative forces are used to integrate the equations of motion, and the conservative forces are applied every $M$ steps as a correction. This reduces the theoretical overhead of a conservative trajectory from a factor of $F\approx 2$ to one of $1+(F-1)/M$. The results using this technique in Table 2, Figure 2 and in Appendix I are essentially indistinguishable from fully conservative ones, using $M=8$, which leads to a small, approximately $20$% slowdown compared to a direct-force, non-conservative trajectory. Appendix I contains further explanation of the MTS technique, as well as more detailed results for MTS simulations using models trained on the water and OC20 datasets. The technique can be also successfully used for constant-pressure simulations, as shown in Section G.3.
@@ Appendix B Theoretical computational cost
However, in the case of backward force evaluation, operations where $x$ is an internal representation and $y$ is a weight can save the $\partial V/\partial y$ calculation. In a simple multi-layer perceptron, where all linear layers correspond to this type of computation, this would yield a backward propagation that is roughly as expensive as the forward pass. This is not the case in transformers, as the attention mechanism involves comparatively expensive operations where both $x$ and $y$ are internal representations, and one can expect the backward propagation of gradients to be somewhere between 1×1$\times$1 × and 2×2$\times$2 × the cost of the forward evaluation.
@@ G.2 NVT
ORB (NC) 300.4${}_{\text{(0.1)}}$ 304.2${}_{\text{(0.2)}}$ 351.0${}_{\text{(0.6)}}$ 301.0${}_{\text{(0.1)}}$ 336.2${}_{\text{(0.8)}}$ 229.5${}_{\text{(1.6)}}$
MACE (C) 300.1${}_{\text{(0.1)}}$ 300.4${}_{\text{(0.2)}}$ 300.1${}_{\text{(0.4)}}$ 300.1${}_{\text{(0.1)}}$ 299.0${}_{\text{(0.6)}}$ 302.3(1.${}_{\text{(1.1)}}$
SevenNet (C) 300.0${}_{\text{(0.1)}}$ 300.1${}_{\text{(0.2)}}$ 299.9${}_{\text{(0.5)}}$ 300.1${}_{\text{(0.1)}}$ 299.9${}_{\text{(0.4)}}$ 300.5(1.${}_{\text{(1.1)}}$
PET-NC (NC) 300.1${}_{\text{(0.1)}}$ 301.4${}_{\text{(0.2)}}$ 312.8${}_{\text{(0.5)}}$ 300.3${}_{\text{(0.1)}}$ 295.6${}_{\text{(0.3)}}$ 309.9${}_{\text{(0.6)}}$
PET-C (C) 300.1${}_{\text{(0.1)}}$ 300.1${}_{\text{(0.2)}}$ 300.3${}_{\text{(0.7)}}$ 300.1${}_{\text{(0.1)}}$ 299.6${}_{\text{(0.3)}}$ 301.0${}_{\text{(0.7)}}$
SOAP-BPNN-NC (NC) 301.9${}_{\text{(0.1)}}$ 340.7${}_{\text{(0.2)}}$ $1.2\cdot 10^{6}$ 309.7${}_{\text{(0.1)}}$ 265.5${}_{\text{(0.2)}}$ 399.5${}_{\text{(0.4)}}$
SOAP-BPNN-C (C) 299.9${}_{\text{(0.1)}}$ 299.8${}_{\text{(0.4)}}$ 300.5${}_{\text{(0.5)}}$ 300.1${}_{\text{(0.2)}}$ 298.9${}_{\text{(0.7)}}$ 302.4${}_{\text{(1.5)}}$
@@ G.6 Generalized Langevin equations
ORB (NC) 1181.5${}_{\text{(5.2)}}$ 1099.2${}_{\text{(4.6)}}$ 1362.5(7.${}_{\text{(7.7)}}$
MACE (C) 303.1${}_{\text{(1.2)}}$ 301.7(1.${}_{\text{(1.1)}}$ 306.0${}_{\text{(1.6)}}$
SevenNet (C) 304.1${}_{\text{(0.8)}}$ 302.6${}_{\text{(1.3)}}$ 307.3${}_{\text{(2.4)}}$
PET-NC (NC) 524.1${}_{\text{(4.6)}}$ 495.5${}_{\text{(4.0)}}$ 585.5${}_{\text{(5.9)}}$
PET-C (C) 301.8${}_{\text{(1.2)}}$ 301.6${}_{\text{(1.2)}}$ 302.2${}_{\text{(1.6)}}$
@@ Appendix I Multiple time stepping
where $\mathbf{q}$ and $\mathbf{p}$ indicate the vectors containing positions and momenta of all particles, and $\mathbf{m}$ the masses. Forces have to be recomputed if they are needed after a position update. In the limit of ideal time scale separation, the resulting dynamics (and the sampled ensemble for constant-temperature simulations) is consistent with $V_{\text{slow}}$, at a much-reduced cost: Letting $F$ denote the slowdown of “slow” over “fast” forces, and assuming that an evaluation of “slow” forces also yields the “fast” ones, MTS reduces the theoretical cost of well-behaved molecular dynamics from $F$ to $1+(F-1)/M$. Choosing $M$ therefore requires a trade-off between accuracy and speed. Morrone et al. (2011) discusses the failure modes of MTS at high $M$, and shows how to use thermostats for mitigation.
@@
ORB, MACE-MP-0, Equiformer, SevenNet calculations were performed using the publicly available repositories, respectively https://github.com/orbital-materials/orb-models/ (orb-models 0.4.0), https://github.com/ACEsuit/mace (mace-torch 0.3.8), https://github.com/FAIR-Chem/fairchem (fairchem-core 1.3.0), and https://github.com/MDIL-SNU/SevenNet (sevenn 0.10.1).